OpenStudy (anonymous):

xSquared - 2x -13 = 0 using the india method for solving quadratic equations

OpenStudy (campbell_st):

its a method that uses complete the square start by adding 13 to both sides of the equation. \[x^2 - 2x = 13\] the coefficient of x^2 is 1 so multiply every term in the quadratic by 4 times the coefficient of x^2 in your question its 4 x 1 = 4 so multiply every term by 4 gives \[4x^2 - 8x = 52\] now go back to the original equation and square the coefficient of x \[(-2)^2 = 4\] now add this to both sides of the equation and you get \[4x^2 - 8x + 4 = 52 + 4\] which means you have \[4x^2 - 8x + 4 = 56\] the left hand side is a perfect square \[(2x - 2)^2 = 56\] you should be able to find the solution from here. here is a link with notes and an example http://www.algebra.com/algebra/homework/quadratic/THEO-2011-08-28-02.lesson